# Two-Step Regularized HARX to Measure Volatility Spillovers in Multi-Dimensional Systems

### v2

@@ Abstract
We identify volatility spillovers across commodities, equities, and treasuries using a hybrid HAR-ElasticNet framework on daily realized volatility for six futures markets over 2002–2025. Our two step procedure estimates own-volatility dynamics via OLS to preserve persistence (roughly 0.99), then applies ElasticNet regularization to cross-market spillovers. The sparse network structure that emerges shows equity markets (ES, NQ) act as the primary volatility transmitters, while crude oil (CL) ends up being the largest receiver of cross-market shocks. Agricultural commodities stay isolated from the larger network. A simple univariate HAR model achieves equally performing point forecasts as our model, but our approach reveals network structure that univariate models cannot. Joint Impulse Response Functions trace how shocks propagate through the network. Our contribution is to demonstrate that hybrid estimation methods can identify meaningful spillover pathways while preserving forecast performance.
@@ 1 Introduction
Think back to early 2020. Equity markets were crashing, crude oil prices were swinging wildly, and grain traders were watching their screens nervously. Soybean futures sometimes seem to catch volatility from crude oil and other ’outside’ markets. This time, however, soybeans futures remained fairly calm compared to what was going on in other markets. One of the most interesting features of financial markets in modern times is the potential for cross-market volatility spillovers among markets that have merely a tangential relationship. Understanding which cross-market linkages actually matter (and which are just noise) is essential for risk management, portfolio construction, and systemic risk monitoring. The problem? Existing approaches face a fundamental tension. Unrestricted multivariate models overfit, but univariate models ignore cross-market dynamics entirely.
@@
The volatility spillover literature has been dominated by GARCH-based methods since Engle (1982) and Bollerslev (1986) established the foundational framework. Extensions to capture asymmetric responses to positive and negative shocks include the GJR-GARCH model of Glosten et al. (1993), which remains widely used for modeling the leverage effect in equity volatility. For multivariate settings, Engle (2002) introduced DCC-GARCH, which models time-varying correlations parsimoniously, while Engle and Kroner (1995) developed the BEKK parameterization that allows explicit modeling of volatility transmission across assets. These multivariate GARCH specifications remain the dominant way to study volatility spillovers—appearing in hundreds of empirical applications across equity, commodity, and currency markets.
@@
. A parallel literature has developed around the spillover index of Diebold and Yilmaz (2009) and Diebold and Yilmaz (2012), which uses variance decompositions from vector autoregressions to measure connectedness across markets. This approach has an advantage. It provides intuitive measures of directional spillovers (how much of asset A’s forecast error variance is attributable to shocks from asset B) without imposing the parametric structure of GARCH. Baruňík and Křehlík (2018) extended this framework to the frequency domain, decomposing spillovers into short-run and long-run components. However, many Diebold-Yilmaz applications use squared returns rather than range-based or high-frequency realized volatility estimators. And the underlying VARs? Usually unrestricted, which becomes problematic as the number of assets grows.
@@
For realized volatility forecasting specifically, Corsi (2009) developed the Heterogeneous Autoregressive (HAR) model, capturing the long-memory properties of volatility using a simple additive structure of daily, weekly, and monthly volatility components. The HAR model has become a workhorse for univariate volatility forecasting. But it doesn’t address cross-market spillovers. Extensions incorporating exogenous variables, like the HAR-X framework, have been developed by Andersen et al. (2007) for jump components and applied to volatility transmission by Bubák et al. (2011) and Corsi and Renò (2012). Audrino and Knaus (2016) demonstrate that LASSO regularization effectively identifies sparse coefficient structures in HAR models when many potential predictors are available. However, these extensions typically focus on forecasting rather than impulse response analysis.
@@
A growing literature applies regularization methods to HAR and VAR models for spillover analysis. Demirer et al. (2018) demonstrated the use of LASSO-VAR within the Diebold and Yilmaz (2012) spillover framework. With this approach they were able to estimate the high-dimensional network of 150 global banks. For commodity markets, Yang and Zhou (2021) apply LASSO-VAR to 25 commodity futures, finding that energy commodities serve as primary spillover transmitters. In the HAR literature, Ma et al. (2018) show that Elastic Net and LASSO outperform individual HAR models for oil price volatility—with Elastic Net performing better due to its ability to handle correlated predictors. Audrino and Knaus (2016) provide theoretical and empirical support for applying LASSO to HAR models, showing that regularization effectively recovers sparse structures without sacrificing forecast accuracy. Ding et al. (2021) find that multivariate LASSO models outperform univariate HAR at longer forecast horizons and introduce ordered LASSO to respect time series structure. An important finding from this literature? Regularization enables network identification even when it doesn’t substantially improve forecasting. Buccheri et al. (2021) document that cross-market spillovers in multivariate HAR models have long-term forecasting power, but LASSO lag selection doesn’t outperform parsimonious HAR specifications. That result foreshadows our own finding: HAR slightly outperforms our multivariate models while the multivariate structure remains essential for understanding spillover dynamics.
@@
Our approach bridges these literatures by developing a hybrid HAR-ElasticNet framework that addresses two distinct econometric challenges in high-dimensional volatility modeling: capturing realistic persistence dynamics and identifying sparse cross-market spillover networks. Standard unrestricted VAR estimation becomes infeasible when the number of variables approaches the sample size. With six markets and three HAR horizons per market, our model includes 108 potential parameters. Our two-step procedure first estimates univariate Heterogeneous Autoregressive (HAR) models via OLS for each market’s own-volatility dynamics, preserving the characteristic high persistence ($\approx 0.99$ at daily frequency) documented in the realized volatility literature (Corsi, 2009). We then apply ElasticNet regularization to the cross-market terms, setting approximately 70–75 percent of spillover parameters to zero. This sparsity is economically sensible—while a handful of central markets exhibit significant cross-linkages, most market pairs have negligible direct volatility transmission. By separating own-dynamics from cross-market selection, our approach combines the strengths of HAR models (realistic persistence) with regularization methods (sparse network identification), yielding impulse responses that reveal volatility spillover structure.
@@
Second, consistent with Buccheri et al. (2021), the parsimonious HAR model and our hybrid HAR-ElasticNet achieve identical out-of-sample RMSE (0.0044). This equivalence reflects the hybrid model’s sparsity. For forecasting purposes, the hybrid model performs similarly to the univariate HAR. This demonstrates that cross-market information contributes negligibly to predictive accuracy while revealing economically meaningful network structure.
Third, Joint Impulse Response Functions (JIRFs) allow us to show how volatility shocks transmit and accumulate at multiple horizons. While the cross effects on our hybrid model were all near zero, the JIRFs allow us to show that in some cases, like equity shocks on on crude oil futures realized volatility, we observed a sharp widening of confidence intervals. This demonstrates how volatility spillovers contribute to ’volatility of volatility’.
Our contribution is both methodological and empirical. We extend the regularized HAR literature (Buccheri et al., 2021; Ma et al., 2018) to a multi-asset-class setting spanning commodities, equities, and treasuries, and combine spillover network identification with Joint IRF analysis for policy-relevant insights into cross-market volatility transmission. To the best of our knowledge, this is the first application of the JIRF concept to a multivariate volatility modeling exercise.
The rest of the paper proceeds as follows. Section 2 describes our data and how we construct realized volatility using the Yang-Zhang estimator. Section 3 lays out the hybrid HAR-ElasticNet model along with the Joint Impulse Response methodology. Section 4 presents results on forecast performance and spillover patterns. Section 5 concludes.
@@ 2.2 Realized Volatility Estimation
The Yang-Zhang estimator is more efficient than Garman-Klass (Garman and Klass, 1980) or Parkinson (Parkinson, 1980) estimators because it’s robust to opening jumps (overnight gaps) while still using all available daily price information. Yang and Zhang (2000) show it has the minimum variance among unbiased range-based estimators for assets that exhibit opening jumps—which is common in futures markets.
@@ 3.3 Hybrid Two-Step Estimation Procedure
The hybrid procedure separates two distinct economic phenomena that warrant different statistical treatments. Own-volatility persistence reflects market-specific information processing—news announcements, supply shocks, and idiosyncratic events generate volatility clustering that persists regardless of cross-market linkages. This component should exhibit high persistence, a property well-captured by HAR’s multi-horizon aggregation and OLS estimation without shrinkage. Cross-market spillovers reflect economic and informational linkages: common macroeconomic exposures, portfolio rebalancing, and contagion during stress episodes. Most market pairs have negligible direct linkages; only a sparse subset of connections transmit volatility materially. ElasticNet regularization identifies this sparse network structure. By fixing own-lag coefficients before estimating spillovers, we ensure that impulse response functions exhibit realistic dynamics: shocks to market $i$ generate persistent own-effects (reflecting HAR decay patterns) combined with sparse cross-transmission (reflecting network structure).
@@ 3.5 Spillover Network Identification
The advantage of the hybrid HAR-ElasticNet framework is that the estimated coefficient matrix directly reveals the spillover network structure. The $\gamma$ coefficients indicate which cross-market transmission channels are economically meaningful (those that survive regularization) versus which are noise that gets shrunk to zero.
@@ 3.6 Joint Impulse Response Functions
Joint Shock Construction. When shocking multiple variables simultaneously (e.g., both ES and NQ in an equity shock), we must specify how the shocks are correlated. Following Wiesen and Beaumont (2024), we use generalized impulse response functions (GIRFs ) in the spirit of Pesaran and Shin (1998), which condition on the estimated residual covariance structure rather than imposing an arbitrary ordering as in Cholesky decomposition. Specifically, when we shock variable $i$ by one standard deviation, we simultaneously adjust other shocked variables $j\in\mathcal{S}$ according to their conditional expectation given the covariance matrix $\hat{\Sigma}_{u}$ of the hybrid HAR-ElasticNet residuals. For the equity shock, this means that the ES and NQ shocks are correlated according to their estimated contemporaneous correlation (approximately 0.85 in our sample), reflecting the empirical reality that equity volatility shocks do not occur independently across indices. This approach avoids the ordering sensitivity of Cholesky-based identification while preserving the joint structure of multi-market volatility events. This approach loses causal interpretation, but implements shocks that are empirically consistent with observed correlations.
Hybrid HAR-ElasticNet IRF Computation. We compute impulse response functions via forward simulation while maintaining the full HAR lag structure. At each horizon $h$, we compute the HAR features (daily, weekly average, monthly average) from the simulated history and apply the estimated hybrid HAR-ElasticNet coefficients to generate the next-period prediction. The IRF is the difference between the shocked and baseline simulation paths. This simulation-based approach naturally preserves the multi-scale transmission dynamics inherent in the HAR framework, as shocks propagate through all three horizon components.
@@ 3.7 Bootstrap Confidence Intervals
    Model Re-estimation: For each bootstrap sample, we re-estimate the hybrid HAR-ElasticNet model. To ensure computational feasibility, we fix the regularization parameters ( $\alpha$, $\lambda$) to their optimal values determined from the full sample cross-validation.
@@ 3.8 Estimation Strategy
We employ the hybrid HAR-ElasticNet model (described in Section 3.3) for both forecast evaluation and network identification. This unified framework preserves realistic own-lag persistence (around 0.99) via OLS before applying ElasticNet regularization to cross-market terms only.
Forecast Evaluation (Section 4.1): To assess out-of-sample predictive accuracy, we estimate the hybrid HAR-ElasticNet model using an 80/20 train-test split. The model is estimated on the first 80% of observations (May 2002 – July 2020, 4,559 observations) and evaluated on the held-out 20% test set (July 2020 – January 2025, 1,140 observations) via rolling one-step-ahead forecasts. This rigorous out-of-sample evaluation prevents overfitting bias and provides unbiased forecast accuracy metrics.
Structural Spillover Analysis (Sections 4.2–4.5): For network identification and JIRF analysis, we estimate the hybrid HAR-ElasticNet model using the full sample (May 2002 – January 2025). Using all available data yields the most accurate measurement of spillover dynamics and volatility transmission mechanisms. Statistical uncertainty is quantified via block bootstrap with model re-estimation (1,000 samples).
@@ 4 Empirical Results
Our hybrid HAR-ElasticNet model separates own-volatility dynamics (estimated via OLS with no regularization) from cross-market spillovers (estimated via ElasticNet). For the cross-market ElasticNet step, the regularization parameter $\alpha$ was set via 5-fold cross validation, and we use $\lambda=0.5$ to balance L1 (sparsity) and L2 (grouping) penalties. This two-step procedure preserves own-lag persistence (around 0.99 for all assets) but still identifies sparse cross-market transmission. See Table 2.
@@ 4.1 Forecast Performance
Consistent with Buccheri et al. (2021), parsimonious specifications without cross-market spillovers achieve equivalent out-of-sample forecast performance to multivariate models. Table 3 reports out-of-sample performance using an 80/20 train-test split: models are estimated on the first 4,559 observations (May 2002 to July 2020) and evaluated on the remaining 1,140 observations (July 2020 to January 2025) via rolling one-step-ahead forecasts. [^2] See the online supplement for MAE and MAPE estimates of forecast performance comparisons. They demonstrate identical results for both models. Both the univariate HAR benchmark and the hybrid HAR-ElasticNet achieve identical average RMSE of 0.0044. The hybrid model’s extreme sparsity—only 8% of cross-market coefficients are nonzero—effectively reduces it to a univariate HAR specification for forecasting purposes, confirming that cross-market spillovers contribute negligibly to predictive accuracy, despite their economic significance for network structure (Sections 4.2–4.5).
This equivalence between forecast accuracy and model complexity is central to the spillover literature. The HAR model exploits the well-documented long-memory properties of realized volatility through its three-component autoregressive structure (Corsi, 2009), efficiently capturing own-market persistence that dominates any predictive signal from cross-market dynamics. The hybrid HAR-ElasticNet model’s aggressive regularization on cross-market terms yields identical forecasts to univariate HAR while preserving the ability to identify economically meaningful spillover channels. This demonstrates that volatility transmission networks, though critical for understanding systemic risk propagation, provide minimal incremental forecasting power beyond own-market persistence.
@@
Figure 2 illustrates the forecast performance visually by plotting actual realized volatility against model predictions for the final 200 days of the test period (approximately September 2024 to January 2025). Both models track the general volatility patterns nearly identically, with HAR (orange dotted) and hybrid HAR-ElasticNet (blue dashed) overlapping almost perfectly. This is consistent with the identical RMSE metrics in Table 3. The deviations from actual values are most noticeable during volatility spikes, where both models underpredict peak realized volatility. This is a well-documented feature of volatility forecasting models that reflects the difficulty of predicting extreme events. The key observation? The forecast errors are small in economic magnitude: typical deviations are 1-2 percentage points of annualized volatility, confirming that both models provide reliable forecasts for risk management purposes. The hybrid HAR-ElasticNet framework (Section 3.3) preserves realistic own-market persistence while identifying sparse spillover structure (Sections 4.2–4.5) without sacrificing forecast accuracy.
Figure 2: Out-of-Sample Volatility Forecasts: Actual vs Model Predictions. Notes: Final 200 days of test period (September 2024 to January 2025). Black solid line shows actual Yang-Zhang realized volatility. Blue dashed line shows hybrid HAR-ElasticNet one-step-ahead forecasts. Orange dotted line shows univariate HAR forecasts. Both models track realized volatility nearly identically with typical forecast errors of 1-2 percentage points, achieving identical RMSE of 0.0044 (Table 3). The hybrid model’s extreme sparsity (8% nonzero cross-market coefficients) effectively reduces it to univariate HAR for forecasting purposes.

### v3

@@ Abstract
We identify volatility spillovers across commodities, equities, and treasuries using a hybrid HARX-ElasticNet framework on daily realized volatility for six futures markets over 2002–2025. Our two step procedure estimates own-volatility dynamics via OLS to preserve persistence, then applies ElasticNet regularization to cross-market spillovers. The sparse network structure that emerges shows equity markets (ES, NQ) act as the primary volatility transmitters, while crude oil (CL) ends up being the largest receiver of cross-market shocks. Agricultural commodities stay isolated from the larger network. A simple univariate HAR model achieves equally performing point forecasts as our model, but our approach reveals network structure that univariate models cannot. Joint Impulse Response Functions trace how shocks propagate through the network. Our contribution is to demonstrate that hybrid estimation methods can identify meaningful spillover pathways while preserving forecast performance.
@@ 1 Introduction
During the market chaos of March 2020 equity markets were crashing, crude oil prices were swinging wildly, and grain traders were watching their screens nervously. Soybean futures sometimes seem to catch volatility from crude oil and other ’outside’ markets. This time, however, soybean futures remained fairly calm compared to what was going on in other markets. One of the most interesting features of financial markets in modern times is the potential for cross-market volatility spillovers among markets that have merely a tangential relationship. Understanding which cross-market linkages actually matter (and which are just noise) is essential for risk management, portfolio construction, and systemic risk monitoring. The problem? Existing approaches face a fundamental tension. Unrestricted multivariate models overfit, but univariate models ignore cross-market dynamics entirely.
@@
The volatility spillover literature has been dominated by GARCH-based methods since Engle (1982) and Bollerslev (1986) established the foundational framework. Extensions to capture asymmetric responses to positive and negative shocks include the GJR-GARCH model of Glosten et al. (1993), which remains widely used for modeling the leverage effect in equity volatility. For multivariate settings, Engle (2002) introduced DCC-GARCH, which models time-varying correlations parsimoniously, while Engle and Kroner (1995) developed the BEKK parameterization that allows explicit modeling of volatility transmission across assets. These multivariate GARCH specifications remain the dominant way to study volatility spillovers—appearing in thousands of empirical applications across many market contexts.
@@
A parallel literature has developed around the spillover index of Diebold and Yilmaz (2009) and spillover shares of Diebold and Yilmaz (2012), which uses variance decompositions from vector autoregressions to measure connectedness across markets. This approach has an advantage. It provides intuitive measures of directional spillovers (how much of asset A’s forecast error variance is attributable to shocks from asset B) without imposing the parametric structure of GARCH. Baruňík and Křehlík (2018) extended this framework to the frequency domain, decomposing spillovers into short-run and long-run components.
@@
For realized volatility forecasting specifically, Corsi (2009) developed the Heterogeneous Autoregressive (HAR) model, capturing the long-memory properties of volatility using a simple additive structure of daily, weekly, and monthly volatility components. The HAR model has become a workhorse for univariate volatility forecasting. But it doesn’t address cross-market spillovers. Extensions incorporating exogenous variables, like the HAR-X framework, have been developed by Degiannakis and Filis (2017), and Audrino and Knaus (2016) demonstrate that LASSO regularization effectively identifies sparse coefficient structures implied by the HAR model. However, these extensions typically focus on forecasting rather than impulse response analysis.
@@
A growing literature applies regularization methods to HAR and VAR models for spillover analysis. Demirer et al. (2018) demonstrated the use of LASSO-VAR within the Diebold and Yilmaz (2012) spillover framework. With this approach they were able to estimate the high-dimensional network of 150 global banks. For commodity markets, Yang et al. (2021) apply LASSO-VAR to 25 commodity futures, finding that energy commodities serve as primary spillover transmitters. Ma et al. (2018) show that ElasticNet and LASSO combinations outperform the oil price volatility benchmark. Audrino and Knaus (2016) provide theoretical and empirical support for applying LASSO to HAR models, showing that regularization effectively recovers sparse structures without sacrificing forecast accuracy. Ding et al. (2021) find that multivariate LASSO models outperform univariate HAR at longer forecast horizons and introduce ordered LASSO to respect time series structure.
@@
Our approach bridges these literatures by developing a hybrid HARX-ElasticNet framework that addresses two distinct econometric challenges in high-dimensional volatility modeling: capturing realistic persistence dynamics and identifying sparse cross-market spillover networks. Standard unrestricted VAR estimation becomes infeasible when the number of variables approaches the sample size. With six markets and three HAR horizons per market, our model includes 108 potential parameters. Our two-step procedure first estimates univariate Heterogeneous Autoregressive (HAR) models via OLS for each market’s own-volatility dynamics, preserving the characteristic high persistence documented in the realized volatility literature (Corsi, 2009). We then apply ElasticNet regularization to the cross-market terms, which set 70–75 percent of spillover parameters to zero. This sparsity is economically sensible—while a handful of central markets exhibit significant cross-linkages, most market pairs have negligible direct volatility transmission. By separating own-dynamics from cross-market selection, our approach combines the strengths of HAR models (realistic persistence) with regularization methods (sparse network identification), yielding impulse responses that reveal volatility spillover structure.
@@
Second, the parsimonious HAR model and our hybrid HARX-ElasticNet achieve effectively identical out-of-sample RMSE (0.0044). This equivalence reflects the hybrid model’s sparsity. For forecasting purposes, the hybrid model performs similarly to the univariate HAR. This demonstrates that cross-market information contributes negligibly to predictive accuracy while revealing economically meaningful network structure.
Third, Joint Impulse Response Functions (JIRFs) allow us to show how volatility shocks transmit and accumulate at multiple horizons. While the cross effects on our hybrid model were all near zero, the JIRFs allow us to show that in some cases, like equity shocks on crude oil futures realized volatility, we observed a sharp widening of confidence intervals. This demonstrates how volatility spillovers contribute to ’volatility of volatility’.
Our contribution is both methodological and empirical. We extend the regularized HAR literature to a multi-asset-class setting spanning commodities, equities, and treasuries, and combine spillover network identification with Joint IRF analysis for policy-relevant insights into cross-market volatility transmission. To the best of our knowledge, this is the first application of the JIRF concept to a multivariate volatility modeling exercise.
The rest of the paper proceeds as follows. Section 2 describes our data and how we construct realized volatility using the Yang-Zhang estimator. Section 3 lays out the hybrid HARX-ElasticNet model along with the Joint Impulse Response methodology. Section 4 presents results on forecast performance and spillover patterns. Section 5 concludes.
@@ 2.2 Realized Volatility Estimation
The Yang-Zhang estimator is more efficient than Garman-Klass (Garman and Klass, 1980) or Parkinson (Parkinson, 1980) estimators because it is robust to opening jumps (overnight gaps) while still using all available daily price information. Yang and Zhang (2000) show it has the minimum variance among unbiased range-based estimators for assets that exhibit opening jumps—which is common in futures markets.
@@ 3.3 Hybrid Two-Step Estimation Procedure
The hybrid procedure separates two distinct economic phenomena that warrant different statistical treatments. Own-volatility persistence reflects market-specific information processing. News announcements, supply shocks, and idiosyncratic events generate volatility clustering that persists regardless of cross-market linkages. This component should exhibit high persistence, a property well-captured by HAR’s multi-horizon aggregation and OLS estimation without shrinkage. Cross-market spillovers reflect economic and informational linkages: common macroeconomic exposures, portfolio rebalancing, and contagion during stress episodes. Most market pairs have negligible direct linkages; only a sparse subset of connections transmit volatility materially. ElasticNet regularization identifies this sparse network structure. By fixing own-lag coefficients before estimating spillovers, we ensure that impulse response functions exhibit realistic dynamics: shocks to market $i$ generate persistent own-effects (reflecting HAR decay patterns) combined with sparse cross-transmission (reflecting network structure).
@@ 3.5 Spillover Network Identification
The advantage of the hybrid HARX-ElasticNet framework is that the estimated coefficient matrix directly reveals the spillover network structure. The $\gamma$ coefficients indicate which cross-market transmission channels are economically meaningful (those that survive regularization) versus which are noise that gets shrunk to zero.
@@ 3.6 Joint Impulse Response Functions
Joint Shock Construction. When shocking multiple variables simultaneously (e.g., both ES and NQ in an equity shock), we must specify how the shocks are correlated. Following Wiesen and Beaumont (2024), we use generalized impulse response functions (GIRFs ) in the spirit of Pesaran and Shin (1998), which condition on the estimated residual covariance structure rather than imposing an arbitrary ordering as in Cholesky decomposition. Specifically, when we shock variable $i$ by one standard deviation, we simultaneously adjust other shocked variables $j\in\mathcal{S}$ according to their conditional expectation given the covariance matrix $\hat{\Sigma}_{u}$ of the hybrid HARX-ElasticNet residuals. For the equity shock, this means that the ES and NQ shocks are correlated according to their estimated contemporaneous correlation (approximately 0.85 in our sample), reflecting the empirical reality that equity volatility shocks do not occur independently across indices. This approach avoids the ordering sensitivity of Cholesky-based identification while preserving the joint structure of multi-market volatility events. This approach loses causal interpretation, but implements shocks that are empirically consistent with observed correlations.
Hybrid HARX-ElasticNet IRF Computation. We compute impulse response functions via forward simulation while maintaining the full HAR lag structure. At each horizon $h$, we compute the HAR features (daily, weekly average, monthly average) from the simulated history and apply the estimated hybrid HARX-ElasticNet coefficients to generate the next-period prediction. The IRF is the difference between the shocked and baseline simulation paths. This simulation-based approach naturally preserves the multi-scale transmission dynamics inherent in the HAR framework, as shocks propagate through all three horizon components.
@@ 3.7 Bootstrap Confidence Intervals
    Model Re-estimation: For each bootstrap sample, we re-estimate the hybrid HARX-ElasticNet model. To ensure computational feasibility, we fix the regularization parameters ( $\alpha$, $\lambda$) to their optimal values determined from the full sample cross-validation.
@@ 3.8 Estimation Strategy
We employ the hybrid HARX-ElasticNet model (described in Section 3.3) for both forecast evaluation and network identification. This unified framework preserves realistic own-lag persistence (around 0.99) via OLS before applying ElasticNet regularization to cross-market terms only.
Forecast Evaluation (Section 4.1): To assess out-of-sample predictive accuracy, we estimate the hybrid HARX-ElasticNet model using an 80/20 train-test split. The model is estimated on the first 80% of observations (May 2002 – July 2020, 4,559 observations) and evaluated on the held-out 20% test set (July 2020 – January 2025, 1,140 observations) via rolling one-step-ahead forecasts. This rigorous out-of-sample evaluation prevents overfitting bias and provides unbiased forecast accuracy metrics.
Structural Spillover Analysis (Sections 4.2–4.5): For network identification and JIRF analysis, we estimate the hybrid HARX-ElasticNet model using the full sample (May 2002 – January 2025). Using all available data yields the most accurate measurement of spillover dynamics and volatility transmission mechanisms. Statistical uncertainty is quantified via block bootstrap with model re-estimation (1,000 samples).
3.9 AI Assistance Statement
We used generative AI (Claude, Anthropic) as a productivity tool for data cleaning, analysis scripting, and initial text drafting. We followed a workflow where AI-generated portions of the text were systematically revised. All content was reviewed, edited, and verified. All remaining errors are our own.
@@ 4 Empirical Results
Our hybrid HARX-ElasticNet model separates own-volatility dynamics (estimated via OLS with no regularization) from cross-market spillovers (estimated via ElasticNet). For the cross-market ElasticNet step, the regularization parameter $\alpha$ was set via 5-fold cross-validation, and we use $\lambda=0.5$ to balance L1 (sparsity) and L2 (grouping) penalties. This two-step procedure preserves own-lag persistence (around 0.99 for all assets) but still identifies sparse cross-market transmission. See Table 2.
@@ 4.1 Forecast Performance
Consistent with Buccheri et al. (2021), parsimonious specifications without cross-market spillovers achieve equivalent out-of-sample forecast performance to multivariate models. Table 3 reports out-of-sample performance using an 80/20 train-test split: models are estimated on the first 4,559 observations (May 2002 to July 2020) and evaluated on the remaining 1,140 observations (July 2020 to January 2025) via rolling one-step-ahead forecasts. [^2] See the online supplement for MAE and MAPE estimates of forecast performance comparisons. They demonstrate identical results for both models. Both the univariate HAR benchmark and the hybrid HARX-ElasticNet achieve identical average RMSE of 0.0044. The hybrid model’s extreme sparsity—only 8% of cross-market coefficients are nonzero—effectively reduces it to a univariate HAR specification for forecasting purposes, confirming that cross-market spillovers contribute negligibly to predictive accuracy, despite their economic significance for network structure (Sections 4.2–4.5).
This equivalence between forecast accuracy and model complexity is central to the spillover literature. The HAR model exploits the well-documented long-memory properties of realized volatility through its three-component autoregressive structure (Corsi, 2009), efficiently capturing own-market persistence that dominates any predictive signal from cross-market dynamics. The hybrid HARX-ElasticNet model’s aggressive regularization on cross-market terms yields identical forecasts to univariate HAR while preserving the ability to identify economically meaningful spillover channels. This demonstrates that volatility transmission networks, though critical for understanding systemic risk propagation, provide minimal incremental forecasting power beyond own-market persistence.
@@
Figure 2 illustrates the forecast performance visually by plotting actual realized volatility against model predictions for the final 200 days of the test period (approximately September 2024 to January 2025). Both models track the general volatility patterns nearly identically, with HAR (orange dotted) and hybrid HARX-ElasticNet (blue dashed) overlapping almost perfectly. This is consistent with the identical RMSE metrics in Table 3. The deviations from actual values are most noticeable during volatility spikes, where both models underpredict peak realized volatility. This is a well-documented feature of volatility forecasting models that reflects the difficulty of predicting extreme events. The key observation? The forecast errors are small in economic magnitude: typical deviations are 1-2 percentage points of annualized volatility, confirming that both models provide reliable forecasts for risk management purposes. The hybrid HARX-ElasticNet framework (Section 3.3) preserves realistic own-market persistence while identifying sparse spillover structure (Sections 4.2–4.5) without sacrificing forecast accuracy.
Figure 2: Out-of-Sample Volatility Forecasts: Actual vs Model Predictions. Notes: Final 200 days of test period (September 2024 to January 2025). Black solid line shows actual Yang-Zhang realized volatility. Blue dashed line shows hybrid HARX-ElasticNet one-step-ahead forecasts. Orange dotted line shows univariate HAR forecasts. Both models track realized volatility nearly identically with typical forecast errors of 1-2 percentage points, achieving identical RMSE of 0.0044 (Table 3). The hybrid model’s extreme sparsity (8% nonzero cross-market coefficients) effectively reduces it to univariate HAR for forecasting purposes.